OpenStudy (anonymous):

600/3000 please help

OpenStudy (anonymous):

600/3000= (6*100)/(30*100) = 6/ 30 6/30 = ????

OpenStudy (anonymous):

i still dont understand or know the answer

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

6*5 = 30 thus divide both numerator and denominator by 6 (6/6)/(30/6) = 1/5

TheSmartOne (thesmartone):